OpenStudy (anonymous):

When looking for the LCM of this x^2-10,x^2+10x. Would the answer be? x(x-10)(x+10)

Parth (parthkohli):

\[\rm (x^2 - 10)(x^2 + 10x) = x^4 +10x^3- 10x^2 -100x = x(x^3 +10x^2 - 10x-100)\]